OpenStudy (annie96):

Given the liner equation x+y=5, find the slope of its graph. -5 -1 1 5 Given the liner equation 4y=8x-3, find the y-intercept. -3/4 -2 3/4 2

OpenStudy (nurali):

y=mx+b x+y=5 y=-x+5 m=-1 4y=8x-3 y=8x-3/4 y=2x-3/4 m=2

OpenStudy (anonymous):

1) first bring the eqution in to y=mx+c form hear m in the slop... 2) y-intersept means nothing but put x=0 then you will get y value...